# Designed self-assembly of programmable colloidal atom-electron equivalents

### v2

@@ Abstract
To unlock the potential for assembling complex colloidal “molecules”, we investigate a minimal binary system of programmable colloidal atom–electron equivalents (PAE-EE), where electron equivalents (EEs) are multivalent linkers with two distinct types of single-stranded DNA (ssDNA) ends complementary to those ssDNAs on binary programmable atom equivalents (PAEs). We derive a statistical mechanical framework for calculating the effective interaction between PAEs mediated by EEs with arbitrary valency, which quantitatively agrees with simulations that explicitly include EEs. Our analysis reveals an anomalous dependence of PAE-PAE interactions on the EE valency, showing that EE-mediated interactions converge at the large valency limit. Moreover, we identify an optimal EE valency that maximizes the interaction difference between targeted and non-targeted binding pairs of PAEs. These findings offer design principles for targeted self-assembly in PAE-EE systems.
@@
DNA-based materials enable the self-assembly of complex and programmable structures through the precise molecular recognition capabilities of DNA Jones et al. (2015). By attaching single-stranded DNAs (ssDNAs) to colloidal surfaces, DNA-coated particles facilitate the formation of organized superstructures Mirkin et al. (1996); Alivisatos et al. (1996); Laramy et al. (2020), where the dual designability of DNA-encoded targeted structures and colloidal properties allows for the emulation of complex biological processes, merging biological sciences with material engineering, e.g., leading to applications in drug delivery Mura et al. (2013); Hu et al. (2018), biosensing Taton et al. (2000); Cao et al. (2002); Xia and Ni (2024) and immunotherapy Radovic-Moreno et al. (2015); Wang et al. (2019a).
@@
The linker-mediated scheme introduces additional designability for orthogonally regulating the specificity of pair interactions Biancaniello et al. (2005); Xiong et al. (2009); Zhang et al. (2015); Lowensohn et al. (2019); Xia et al. (2020); Kwon et al. (2022). Moreover, multivalent linkers, or termed electron equivalents (EEs), in these systems exhibit behavior analogous to electrons in metals, roaming through colloidal crystals and fostering effective attraction between larger DNA-coated colloids, termed programmable atom equivalents (PAEs) Girard et al. (2019). This system serves as a colloidal analog to atoms and electrons, manifesting phenomena such as EE localization similar to the metal-insulator transition Girard et al. (2019); Lin and Olvera De La Cruz (2020); Ehlen et al. (2021); Lopez-Rios et al. (2021); Lin and Olvera de la Cruz (2022), the emergence of coordination numbers Cheng et al. (2021); Wang et al. (2022), and the formation of colloidal alloys Wang et al. (2019b) and colloidal superionic conductors Lin and Olvera de la Cruz (2023).
@@
Existing programmable atom-electron equivalent (PAE-EE) systems have primarily utilized EEs that are multivalent linkers with identical ssDNA ends. However, this homogeneity limits the complexity of the resulting assemblies. For more sophisticated structures, such as colloidal “molecules”, it is essential to employ EEs with heterogenous ends to guide programmable assembly. In this Letter, we investigate a minimal binary PAE-EE system, where the EEs possess two distinct types of ends complementary with the ssDNA ends on binary PAEs. We derive an explicit formula for calculating the effective interaction between PAEs mediated by EEs of arbitrary valency from a statistical mechanical model based on the quantitative model for DNA-coated colloids Angioletti-Uberti et al. (2013, 2014); Di Michele et al. (2016); Xia et al. (2020). This approach captures the interplay among binding energy, configurational entropy, and combinatorial entropy, allowing us to study the distribution of EEs in different binding states, which serves as a framework for Monte Carlo (MC) simulations with implicit EEs. We study the EE-mediated pair interaction between the same and different types of PAEs and corresponding phase behaviors, offering a design principle to guide the targeted self-assembly in PAE-EE systems.
@@
Model and EE-mediated interactions.
We consider an $\alpha\gamma$-type binary PAE system in a solution containing free EEs, and each PAE is modeled as a hard sphere of radius $R$ coated with $n_{\rm\alpha}$ $\alpha$-type or $n_{\rm\gamma}$ $\gamma$-type mobile ssDNA ends (Fig. 1(a)). The mobility of ssDNAs coated on PAEs allows treating them as a mean-field-like adsorption layer to derive a closed form for effective interactions, and does not change the qualitative physics in the system Angioletti-Uberti et al. (2014). ssDNA ends are modeled as infinitely thin, hard rods of length $l_{r}$. Free EEs bridge PAEs in the solution and are connected to a reservoir of chemical potential $\mu$. Each EE is modeled as a polymer blob of zero radius, which is based on the fact that EEs are normally much smaller than PAEs experimentally. An EE consists of equimolar $\lambda$ number of $\alpha^{\prime}$ and $\gamma^{\prime}$-type ends, which are complementary to the $\alpha$ and $\gamma$-type ssDNAs coated on colloids, respectively. We set the binding energy as $\Delta G_{\rm\alpha\alpha^{\prime}}=\Delta G_{\rm\gamma\gamma^{\prime}}=\Delta G$, which characterizes the chemical details of specific binding. Non-specific bindings between $\alpha-\gamma$, $\alpha-\gamma^{\prime}$, $\alpha^{\prime}-\gamma$, and $\alpha^{\prime}-\gamma^{\prime}$ ends are neglected.
In typical experimental PAE-EE systems, van der Waals interactions and electrostatic interactions can be neglected Varilly et al. (2012). Besides, we also neglect the depletion interaction originating from the EEs acting as depletants, particularly when EEs are sufficiently small. Based on the fact that PAEs are much larger than EEs Girard et al. (2019), we assume that the motion of PAEs and EEs can be decoupled, and EEs move within the static external potential generated by PAEs and provide effective interactions among PAEs. Consequently, PAEs interact mainly through entropic repulsion and EE-mediated attraction:
@@
where $F_{\rm att}$ is the free energy of EEs given the coordinations of PAEs, and $F_{\rm att}^{\prime}$ represents the free energy of EEs for isolated PAE configurations. The entropic repulsion $\beta F_{\rm rep}$ originates from the excluded volume of ssDNAs overlapping with the hard cores of PAEs Angioletti-Uberti et al. (2014). Here $\beta=1/k_{B}T$ with $k_{B}$ and $T$ the Boltzmann constant and the temperature of the system, respectively. As shown in Fig. 1(c), we consider $\beta U_{\rm eff}^{\alpha\gamma/\alpha\alpha}$ to be the targeted/non-targeted PAE-PAE interaction, respectively.
@@
As EEs are much smaller than PAEs, we assume that each EE can only be adsorbed onto one PAE $i$ (adsorbed state $\mathcal{A}_{i}$) or bridge between PAE $i$ and $j$ (bridging state $\mathcal{B}_{ij}$). We note that the EEs here can exhibit different binding substates characterized by the number of bound ends. As shown in Fig. 1(b), we denote $\mathcal{A}_{i}^{k}$ substate as $k$ ends on an EE binding with the ends on particle $i$, and $\mathcal{B}_{ij}^{kl}$ substate as $k$ and $l$ ends on an EE binding with the ends on PAE $i$ and $j$, respectively. Using the saddle-point approximation, we obtain the equilibrium number of EEs, ${m}_{i}^{k}$ and ${q}_{ij}^{kl}$, in different substates, $\mathcal{A}_{i}^{k}$ and $\mathcal{B}_{ij}^{kl}$, respectively (see SM Sec. I sup )
@@
where $\bar{n}_{i}$ is the number of unbound ends on PAE $i$:
$\bar{n}_{i}=n_{i}-\sum_{k}km_{i}^{k}-\sum_{j}\sum_{k,l}kq_{ij}^{kl}.$ (3)
The binding strength of an EE in $\mathcal{A}_{i}^{k}$ substate, $\Xi_{\mathcal{A},i}^{k}$, and in $\mathcal{B}_{ij}^{kl}$ substate, $\Xi_{\mathcal{B},ij}^{kl}$, can be written as
@@
$\Xi_{\mathcal{B},ij}^{kl}$ $=m_{ij}^{0}w^{kl}(\chi_{i})^{k}(\chi_{j})^{l},$ (4b)
where $m_{i}^{0}$ is the number of unbound EEs in the adsorption layer of PAE $i$, while $m_{ij}^{0}$ is the number of unbound EEs in the overlapped adsorption layers between PAE $i$ and $j$. $\chi_{i}$ denotes the individual binding strength of each pair of ends binding between PAE $i$ and an EE, which counts both energy and entropy effects. The combinatorial term in bridging $w^{kl}=\binom{\lambda}{k,l}$ if $t_{i}=t_{j}$, while $w^{kl}=\binom{\lambda}{k}\binom{\lambda}{l}$ if $t_{i}\neq t_{j}$, with $t_{i/j}$ the type of PAE $i/j$. The free energy of bound EEs is given by
@@
where $M_{i}$ and $Q_{ij}$ are total numbers of EEs in the adsorbed state $\mathcal{A}_{i}$ and bridging state $\mathcal{B}_{ij}$, respectively, and the constant indicates the reference to different configurations of PAEs. We notice that Eq. 3 can be numerically solved, and $M_{i}$ and $Q_{ij}$ in Eq. 5 are in closed forms for arbitrary $\lambda$ (see SM Sec. I sup ).
Figure 2: Numerical verification of the effective pair interactions between PAEs. (a,b) EE-mediated effective PAE-PAE interactions $\beta U_{\rm eff}(r)$ as functions of the center-to-center distance $r$ between two PAEs for different EE valency $\lambda$, where the PAEs are of the same type (a, $\beta U_{\rm eff}^{\alpha\alpha}$) and different types (b, $\beta U_{\rm eff}^{\alpha\gamma}$). (c) The ratio of unbound ssDNAs $\bar{n}^{\alpha\gamma}/n$ as a function of $r$ for different $\lambda$ between PAEs of different types. (d) Number of EEs $N_{\rm E}^{\alpha\gamma}$ depicted by the shadow area in the inset as a function of $r$ for different $\lambda$ between PAEs of different types. Symbols are from numerical simulations, and curves are theoretical predictions. Parameters: $\beta\mu=-14$, $\beta\Delta G=-10$, $n=100$ and $l_{r}=0.1R$.
Figure 3: Entropy driven PAE-PAE interactions (a,b) The strength of EE-mediated interaction $\beta U_{\rm eff}(r^{*})$ as functions of individual binding energy $\beta\Delta G$ between two PAEs for different EE valency $\lambda$ at $\beta\mu=-13$, where the PAEs are of the same type (a, $\beta U_{\rm eff}^{\alpha\alpha}(r^{*})$) and different types (b, $\beta U_{\rm eff}^{\alpha\gamma}(r^{*})$). Symbols are from numerical simulations, and curves are theoretical predictions. (c) Singlet fraction in manybody PAE-EEs systems at $\eta=0.094$ as a function of $\beta\Delta G$ for different $\lambda$ at $\beta\mu=-11$. Parameters: $n_{\alpha}=n_{\gamma}=100$, $r^{\ast}=2R$ and $l_{r}=0.1R$.
Numerical verification.

### v3

@@ Abstract
To unlock the potential for assembling complex colloidal “molecules”, we investigate a minimal binary system of programmable colloidal atom–electron equivalents (PAE-EE), where electron equivalents (EEs) are multivalent linkers with two distinct types of single-stranded DNA (ssDNA) ends complementary to those ssDNAs on binary programmable atom equivalents (PAEs). We derive a statistical mechanical framework for calculating the effective interaction between PAEs mediated by EEs with arbitrary valency, which quantitatively agrees with simulations using explicit EEs. Our analysis reveals an anomalous dependence of PAE-PAE interactions on the EE valency, showing that EE-mediated interactions converge at the large valency limit. Moreover, we identify an optimal EE valency that maximizes the interaction difference between targeted and non-targeted binding pairs of PAEs. These findings offer design principles for targeted self-assembly in PAE-EE systems.
: Rep. Prog. Phys.
Keywords: DNA coated colloids, self-assembly, mean-field theory, Monte Carlo simulation
[…]
1 Introduction
@@
DNA-based materials enable the self-assembly of complex and programmable structures through the precise molecular recognition capabilities of DNA [1]. By attaching single-stranded DNAs (ssDNAs) to colloidal surfaces, DNA-coated particles facilitate the formation of organized superstructures [2, 3, 4], where the dual designability of DNA-encoded targeted structures and colloidal properties allows for the emulation of complex biological processes, merging biological sciences with material engineering, e.g., leading to applications in drug delivery [5, 6], biosensing [7, 8, 9] and immunotherapy [10, 11].
@@
The linker-mediated scheme introduces additional designability for orthogonally regulating the specificity of pair interactions [12, 13, 14, 15, 16, 17]. Multivalent linkers, or termed electron equivalents (EEs), in these systems exhibit behavior analogous to electrons in metals, roaming through colloidal crystals and fostering effective attraction between larger DNA-coated colloids, termed programmable atom equivalents (PAEs) [18]. This system serves as a colloidal analog to atoms and electrons, manifesting phenomena such as EE localization similar to the metal-insulator transition [18, 19, 20, 21, 22], the emergence of coordination numbers [23, 24], and the formation of colloidal alloys [25] and colloidal superionic conductors [26].
@@
Existing programmable atom-electron equivalent (PAE-EE) systems have been primarily utilizing EEs that are multivalent linkers with identical ssDNA ends. This homogeneity limits the complexity of the resulting assemblies. For more complex structures, such as colloidal “molecules”, it is essential to employ EEs with heterogeneous ends to guide programmable assembly. Lowensohn et al. [15] and our previous work [16] studied the system of linkers with heterogenous bi-valency, providing a foundation for understanding linker-mediated interactions. However, how the multivalency of EEs influences the phase behavior of the systems remains largely unexplored. In particular, the design rule for selecting the optimal linker valency to achieve the best possible self-assembly remains unknown. Here, we extend the theoretical framework in Refs. [15, 16] and focus on the phenomena that emerge for arbitrary EE valency. We investigate a minimal binary PAE-EE system, where the EEs possess two distinct types of ends complementary to the ssDNA ends on binary PAEs. We derive an explicit formula for calculating the effective interaction between PAEs mediated by EEs of arbitrary valency from a statistical mechanical model of DNA coated colloids [27, 28, 29, 16]. This approach allows us to study the distribution of EEs in different binding states, which can serve as a framework for Monte Carlo (MC) simulations with implicit EEs. We study the EE-mediated pair interaction between the same and different types of PAEs and corresponding phase behaviors, offering a design principle to guide the targeted self-assembly in PAE-EE systems.
2 Methods
2.1 Model
We consider an $\alpha\gamma$-type binary PAE system in a solution containing free EEs, and each PAE is modeled as a hard sphere of radius $R$ coated with $n_{\rm\alpha}$ $\alpha$-type or $n_{\rm\gamma}$ $\gamma$-type mobile ssDNA ends (Fig. 1(a)). The mobility of ssDNAs coated on PAEs allows treating them as a mean-field-like adsorption layer to derive a closed form for effective interactions, and does not change the qualitative physics in the system [28]. ssDNA ends are modeled as infinitely thin, hard rods of length $l_{r}$. Free EEs bridge PAEs in the solution and are connected to a reservoir of chemical potential $\mu$. Each EE is modeled as a polymer blob of zero radius, which is based on the fact that EEs are normally much smaller than PAEs experimentally. An EE consists of equimolar $\lambda$ number of $\alpha^{\prime}$ and $\gamma^{\prime}$-type ends, which are complementary to the $\alpha$ and $\gamma$-type ssDNAs coated on colloids, respectively. We set the binding energy as $\Delta G_{\rm\alpha\alpha^{\prime}}=\Delta G_{\rm\gamma\gamma^{\prime}}=\Delta G$, which characterizes the chemical details of specific binding. Non-specific bindings between $\alpha-\gamma$, $\alpha-\gamma^{\prime}$, $\alpha^{\prime}-\gamma$, and $\alpha^{\prime}-\gamma^{\prime}$ ends are neglected.
@@ 2.1 Model
2.2 EE-mediated interactions
In typical experimental PAE-EE systems, van der Waals interactions and electrostatic interactions can be neglected [30, 31]. Besides, we also neglect the depletion interaction originating from the EEs acting as depletants, particularly when EEs are sufficiently small. Based on the fact that PAEs are much larger than EEs [18], we assume that the motion of PAEs and EEs can be decoupled, and EEs move within the static external potential generated by PAEs and provide effective interactions among PAEs. Consequently, the effective interaction among PAEs can be expressed by
@@ 2.2 EE-mediated interactions
which, when negative, indicates a favorable attractive interaction among the PAEs. $F_{\rm att}$ is the free energy of EEs given the coordinations of PAEs, $F_{\rm att}^{\prime}$ represents the free energy of EEs for isolated PAE configurations, and $\beta=1/k_{B}T$ with $k_{B}$ and $T$ the Boltzmann constant and the temperature of the system, respectively. The entropic repulsion $\beta F_{\rm rep}$ originates from the excluded volume of ssDNAs overlapping with the hard cores of PAEs [28]. We derive a statistical mechanical framework to calculate the effective interaction $\beta F_{\rm att}$ among PAEs mediated by EEs. To make the following derivations more accessible, we briefly outline the physical picture behind our statistical mechanical framework. As EEs are much smaller than PAEs, we assume that each EE can only be adsorbed onto one PAE $i$ (adsorbed state $\mathcal{A}_{i}$) or bridge between PAE $i$ and $j$ (bridging state $\mathcal{B}_{ij}$). In a grand‑canonical description, the solution acts as a reservoir of EEs with a fixed chemical potential $\mu$, and the number of EEs in every possible binding substate is allowed to fluctuate. We derive the corresponding partition function $\mathcal{Z}$ for the bound EEs on PAEs and, by applying the saddle‑point approximation, replace these fluctuating occupations by their most probable values. This yields closed‑form expressions for the average numbers of adsorbed and bridging EEs, $M_{i}$ and $Q_{ij}$. Substituting $M_{i}$ and $Q_{ij}$ into the grand potential leads to an explicit and tractable formula for the EE‑mediated contribution $\beta F_{\rm{att}}$ that can be used directly in analytic calculations or Monte‑Carlo simulations.
@@
We note that EEs here can stay in different binding substates characterized by the number of bound ends. As shown in Fig. 1(b), we denote $\mathcal{A}_{i}^{k}$ substate as $k$ ends on an EE binding with the ends on particle $i$, and $\mathcal{B}_{ij}^{kl}$ substate as $k$ and $l$ ends on an EE binding with the ends on PAE $i$ and $j$, respectively. The grand canonical partition function for the bound EEs on PAEs is (see Supplementary Material Sec. II.A [32]) {strip}
$\mathcal{Z}\left(\{\mathbf{m}_{i},\mathbf{q}_{ij}\}\right)=\sum_{\{\mathbf{m}_{i},\mathbf{q}_{ij}\}}\mathcal{W}\left(\{\mathbf{m}_{i},\mathbf{q}_{ij}\}\right)\prod_{i}\left[\prod_{k}\left({\xi_{\mathcal{A},i}^{k}}\right)^{m_{i}^{k}}\prod_{j<i}\prod_{k,l}\left({\xi_{\mathcal{B},ij}^{kl}}\right)^{q_{ij}^{kl}}\right],$ (2)
where $\mathbf{m}_{i}$ is a tuple of numbers ${m_{i}^{k}}$, where $k=1,\cdots,\lambda$ that represents the number of EE in $\mathcal{A}_{i}^{k}$ substate, and $\mathbf{q}_{ij}$ is a tuple of numbers $q_{ij}^{kl}$ that represents the number of EE in $\mathcal{B}_{ij}^{kl}$ substate. By counting the total number of ends in EE, we note that when $i$ and $j$ are the same types of particles, i.e., $t_{i}=t_{j}$, then for all $k+l\geq\lambda$, $q_{ij}^{kl}=0$. $\{\mathbf{m}_{i},\mathbf{q}_{ij}\}$ accounts for all possible values of $\mathbf{m}_{i}$ and $\mathbf{q}_{ij}$ for all particles and pairs of particles, respectively. The second and third terms represent the energy term for all EEs in adsorbed and bridging states, where $\prod_{i}\cdot$ indicates the product of a sequence on each particle $i$ from 1 to $N$, $\prod_{i,j<i}\cdot$ indicates the product of a sequence on each pair of particles $i$ and $j$, $\prod_{k}=\prod_{k=1}^{\lambda}$ denotes the product on each $\mathcal{A}_{i}^{k}$ adsorbed substate, and $\prod_{k,l}$ denotes the product on each $\mathcal{B}_{ij}^{kl}$ adsorbed substate with
$\prod_{k,l}=\left\{\begin{aligned} \prod_{k=1}^{\lambda-1}\prod_{l=1}^{\lambda-k},\quad t_{i}=t_{j};\\ \prod_{k=1}^{\lambda}\prod_{l=1}^{\lambda},\quad t_{i}\neq t_{j}.\end{aligned}\right.$ (3)
Here, the individual partition function of EEs in substate $\mathcal{A}_{i}^{k}$ and $\mathcal{B}_{ij}^{kl}$ can be expressed as $\xi_{\mathcal{A},i}^{k}=m_{i}^{0}(\chi_{i})^{k}$ and $\xi_{\mathcal{B},ij}^{kl}=m_{ij}^{0}(\chi_{i})^{k}(\chi_{i})^{l}$, respectively. $m_{i}^{0}=V_{\mathcal{A},i}e^{\beta\mu}$ is the number of unbound EEs in the adsorption layer of PAE $i$, while $m_{ij}^{0}=V_{\mathcal{B},ij}e^{\beta\mu}$ is the number of unbound EEs in the overlapped adsorption layers between PAE $i$ and $j$. Here $V_{\mathcal{A},i}$ represents the remaining volume after excluding the hard sphere repulsion from neighboring particles and $V_{\mathcal{B},ij}$ represents the overlapping volume between the two adsorption layers from particle $i$ and $j$ because EE in a bridging state can only stay in this region. $\chi_{i}=e^{-\beta\Delta G/\rho^{\circ}V_{\mathcal{A},i}}$ with $\rho^{\circ}$ the standard concentration denotes the individual binding strength of each pair of ends binding between PAE $i$ and an EE, which accounts for both effects of energy and entropy (see Supplementary Material Sec. II.D [32]). All possible combinations of PAE/EEs end binding can be expressed by (see Supplementary Material Sec. II.E [32]):
$\mathcal{W}\left(\{\mathbf{m}_{i},\mathbf{q}_{ij}\}\right)=$ $\prod_{i}\left[\frac{n_{i}!}{\bar{n}_{i}!}\prod_{k}\frac{\binom{\lambda}{k}^{m_{i}^{k}}}{m_{i}^{k}!}\prod_{j<i}\prod_{k,l}\frac{\left(w^{kl}\right)^{q_{ij}^{kl}}}{q_{ij}^{kl}!}\right],$ (4)
where the combinatorial term in the bridged state
$w^{kl}=\left\{\begin{aligned} \binom{\lambda}{k,l}&=\frac{\lambda!}{k!l!(\lambda-k-l)!},\quad&t_{i}=t_{j};\\ \binom{\lambda}{k}\binom{\lambda}{l}&=\frac{\lambda!\lambda!}{k!l!(\lambda-k)!(\lambda-l)!},\quad&t_{i}\neq t_{j}.\end{aligned}\right.$ (5)
and $\bar{n}_{i}$ denotes the number of unbound ends on particle $i$:
$\bar{n}_{i}=n_{i}-\sum_{k}km_{i}^{k}-\sum_{j}\sum_{k,l}kq_{ij}^{kl}.$ (6)
Using the saddle-point approximation, we obtain the equilibrium number of EEs, ${m}_{i}^{k}$ and ${q}_{ij}^{kl}$, in different substates, $\mathcal{A}_{i}^{k}$ and $\mathcal{B}_{ij}^{kl}$, respectively (see Supplementary Material Sec. II.B [32])
@@
where the binding strength of an EE in $\mathcal{A}_{i}^{k}$ substate, $\Xi_{\mathcal{A},i}^{k}$, and in $\mathcal{B}_{ij}^{kl}$ substate, $\Xi_{\mathcal{B},ij}^{kl}$, can be written as
@@
$\Xi_{\mathcal{B},ij}^{kl}$ $=m_{ij}^{0}w^{kl}(\chi_{i})^{k}(\chi_{j})^{l}.$ (8b)
Thus, the free energy of bound EEs is given by
@@
where $M_{i}$ and $Q_{ij}$ are the total numbers of EEs in the adsorbed state $\mathcal{A}_{i}$ and bridging state $\mathcal{B}_{ij}$, respectively, and the constant indicates the reference to different configurations of PAEs. Similarly, the free energy of EEs for isolated PAE configurations $\beta F_{\rm att}^{\prime}$ is derived in Supplementary Material Sec. II.C. We notice that Eq. 6 can be numerically solved, and $M_{i}$ and $Q_{ij}$ in Eq. 9 are in closed forms for arbitrary $\lambda$ (see Supplementary Material Sec. II.B [32]). With the system consisting of $N$ particles, we obtain a set of $N$-element self-consistent equation with variables $(\bar{n}_{1},\bar{n}_{2},\cdots,\bar{n}_{N})$ from Eq. 6, which can be numerically solved by fixed-point iteration. In Supplementary Material Sec. II.F, we prove that there exists a unique solution $\{\bar{n}_{i}\}$ for $0\leq\bar{n}_{i}\leq n_{i}$ [32]).
The theory is based on the mean field approximation, assuming negligible fluctuations in EE occupancy and rapid DNA binding/unbinding kinetics compared to colloidal dynamics. This approximation holds well when many EEs participate, do not strongly interfere with each other, and binding equilibria can quickly adjust to changes in colloidal configurations. We note that when the EE is heterogeneously bivalent, i.e., $\lambda=1$, the free energy can be simplified as
$\beta F_{\rm att}=\sum_{i}\left(n_{i}\log\frac{\bar{n}_{i}}{n_{i}}+\sum_{j<i}Q_{ij}\right),$ (10)
which is consistent with the bivalent linker-mediated interaction expression [15, 16].
3 Results
3.1 Numerical verification